OpenStudy (anonymous):

George has a small coin collection. He has collected 16 gold coins, 28 silver coins, 32 copper coins, and 4 bronze coins. At a flea market, George sells 15% of his coins. Find the number of coins he sold.

OpenStudy (anonymous):

I will medal and fan!!

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (anonymous):

(number of coins sold)/(total number of coins) X 100

OpenStudy (anonymous):

so how can I do that

OpenStudy (anonymous):

oops, solve for number of coins sold... what i wrote is the percentage of coins sold, but you already know that. is we convert the percent to a number (decimal), what would it be?

OpenStudy (anonymous):

0.15?? @pgpilot326

OpenStudy (anonymous):

yep so (number of coins sold)/(total number of coins) = .15 how many total coins were there?

OpenStudy (anonymous):

16+28+32+4=2852

OpenStudy (anonymous):

? did you multiply?

OpenStudy (anonymous):

No I added all that

OpenStudy (anonymous):

and that's what you got?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

16 + 4 + 32 + 28 = (16+4) + (32+28) = 20 + 60 =?

OpenStudy (anonymous):

what??? tooo confusing

OpenStudy (anonymous):

if you have no sense that adding 4 numbers, all less than 100 must be less than 2852 then you should learn how to add before doing percents...

OpenStudy (anonymous):

how did you add those numbers? what is 16 + 28?

OpenStudy (anonymous):

Just can u tell me the answer I gotta go now

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i'm not here to give out answers unless you are here to pay me

OpenStudy (anonymous):

you here to pay me? each answer, $25

OpenStudy (anonymous):

what